OpenStudy (anonymous):

A rocket, initially at rest on the ground, accelerates straight upward from rest with constant acceleration 44.1m/s^2 . The acceleration period lasts for time 7.00s until the fuel is exhausted. After that, the rocket is in free fall.

OpenStudy (anonymous):

Find the maximum height y_max reached by the rocket. Ignore air resistance and assume a constant acceleration due to gravity equal to 9.80 \rm{m/s^2} .

OpenStudy (anonymous):

y=0.5*(a)*t^2

OpenStudy (anonymous):

propelled flight: y = Vi*t +1/2*a*t^2 Vi=0 y1 = 1/2*44*(7^2) Vf = Vi +at Vi=0 Vf = 44*7 free fall: Vf^2 = Vi^2 +2ad Vf = 0 Vi=44*7 -(44*7)^2 = -2*9.8*y2 y1+y2 = max height